# Neural stochastic Volterra equations: learning path-dependent dynamics

### v2

@@ 1. Introduction
Stochastic Volterra equations (SVEs) are used as mathematical models for the time evolutions of random systems appearing in various areas like biology, finance or physics. SVEs are a natural generalization of ordinary stochastic differential equations (SDEs) and, in contrast to SDEs, they are capable to represent random dynamics with memory effects and very irregular trajectories. For instance, SVEs are used in the modelling of turbulence [BNS08], of volatility on financial markets [EER19] and of DNA patterns [RBS10].
@@ 2.3. Stability for SVEs
Let $\mu,\mu_{n}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma,\sigma_{n}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$, $n\in\mathbb{N}$, be measurable functions such that:
@@
Suppose that $\|K_{\mu}\|_{q}+\|\tilde{K}_{\mu}\|_{q}<\infty$, $\|K_{\sigma}\|_{2\tilde{q}}+\|\tilde{K}_{\sigma}\|_{2\tilde{q}}<\infty$, and $\|g\|_{\infty}+\|\tilde{g}\|_{\infty}<\infty$.
@@ 3.1. Disturbed pendulum equation
Example paths of the training and the testing sets together with their learned approximations are shown in Figure 3.1. It is clearly visible that while DeepONet is not able to generalize properly to the testing set, the learned neural SVE paths are very close to the true paths also for the test set.
{tblr}
—m3cm—m3cm—m3cm—m3cm— Neural SVE:
Training set Neural SVE:
Test set DeepONet:
Training set DeepONet:
@@
The results are presented in Table 2. It stands out immediately that for $n=100$ the performance in the test set is far worse than in the training set. Most solutions remain in the range of $[-\pi,\pi]$. Yet, in some rare cases, solutions may explode outside this range. An example of this can be seen in Figure 3.1. In the training set this happens too rarely for the neural network to learn the functions outside of $[-\pi,\pi]$ precisely. Table 2 clearly shows that this discrepancy disappears when the training set is sufficiently large. In the case of $n=2000$ one also sees that the neural network can learn these complex functions as good as in the prior example. Then the plots for the test set look similar to the training set plots of the Neural SVE trained with $n=100$ trajectories.
@@
{tblr}
—m3cm—m3cm—m3cm—m3cm— Neural SVE:
Training set Neural SVE:
Test set DeepONet:
Training set DeepONet:
@@ 3.2. Rough Heston equation
Example paths of the training and the testing sets together with their learned approximations are shown in Figure 3.2.
{tblr}
—m3cm—m3cm—m3cm—m3cm— Neural SVE:
Training set Neural SVE:
Test set DeepONet:
Training set DeepONet:
@@ 3.3. Generalized Ornstein–Uhlenbeck process
Example paths of the training and the testing sets together with their learned approximations are shown in Figure 3.3.
{tblr}
—m3cm—m3cm—m3cm—m3cm— Neural SVE:
Training set Neural SVE:
Test set DeepONet:
Training set DeepONet:
@@
Example paths of the training and the testing sets together with their learned approximations are shown in Figure 3.3.
{tblr}
—m6cm—m6cm— Neural SVE: Training set Neural SVE: Test set
@@ 3.4. Monetary reserve modelling
Again, it can be observed that the neural SVE learned the dynamics quite well.
@@
In Figure 3.4 one can see the log-monetary reserves of the $10$ banks from one data set and its estimation from the test and the training set each. The dips when the first borrowing contracts end after $t=10$ are not as pronounced in the network’s approximation as in the original data set. This likely stems from neural networks struggling to learn indicator functions such as $K_{\mu}$ here. Still, the order of the banks’ reserves as well as their magnitude are captured very well.
@@ 3.5. Comparison to neural SDEs
where $\xi\sim\mathcal{N}(5,0.5)$ and $T=5$. The process $(X_{t})_{t\in[0,5]}$ is expected to decrease in the first quarter of the interval $[0,5]$ where $K(t-s)=1$ holds due to the mean-reverting effect of the drift coefficient $\mu(s,x)=2-x$, then something unpredictable will happen and finally in the last part of the interval $t\in[0,5]$ where the kernels attain $-1$ for a large proportion of $s\in[0,t]$, the process might become big due to the turning sign in the drift. Hence, it is expected that the path dependency will have a substantial impact.
@@
Example paths of the training and the testing sets together with their learned approximations are shown in Figure 3.5.
{tblr}
—m3cm—m3cm—m3cm—m3cm— Neural SVE:
Training set Neural SVE:
Test set Neural SDE:
Training set Neural SDE:
@@ 3.6. Computational aspects
Using Prokhorov’s theorem and the Skorokhod representation theorem, one deduces that there is a probability space $(\hat{\Omega},\hat{\mathcal{F}},\hat{\mathbb{P}})$ with continuous stochastic processes $\hat{X}^{l},\hat{Y}^{l},\hat{B}^{l},\hat{A}^{l},\hat{M}^{l}$, $l\in\mathbb{N}$ and $\hat{X},\hat{Y},\hat{B},\hat{A},\hat{M}$ such that
@@
$(\hat{X}^{l},\hat{Y}^{l},\hat{B}^{l},\hat{A}^{l},\hat{M}^{l})\to(\hat{X},\hat{Y},\hat{B},\hat{A},\hat{M})$
in $C([0,T];\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\times\mathbb{R}^{d})$ as $l\to\infty$, $\hat{\mathbb{P}}$-a.s., and $\hat{\xi}^{l}\to\hat{\xi}$ as $l\to\infty$, $\hat{\mathbb{P}}$-a.s. [^1] One may drop the $\xi,\hat{\xi}^{l}$, $l\in\mathbb{N}$, here, as they are uniquely determined by the $X$, $\hat{X}^{l}$, respectively. With $\stackrel{{\scriptstyle\mathscr{D}}}{{\sim}}$ we denote equality in law. From here on we identify any space of continuous functions with the supremum norm.
@@
We check that $(\hat{X},\hat{Z})$ with $\hat{Z}:=\hat{A}+\hat{M}$, $(\hat{\Omega},\hat{\mathcal{F}},\hat{\mathbb{P}})$, (^$(\hat{\mathcal{F}}_{t})_{t\in[0,T]}$ solves the Volterra local martingale problem [PS23a, Definition 2.4] given $(\xi g,\mu,\sigma,K_{\mu},K_{\sigma})$. Then, by [PS23a, Lemma 2.7] $\hat{Y}$ is a solution of the SVE (2.1). Conditions (i)-(iii) of [PS23a, Definition 2.4] are clear. (iv) of [PS23a, Definition 2.4] follows as in the proof of [PS23a, Lemma 3.9].
@@
We know that $\hat{Y}^{l}\to\hat{Y}$ and that $\hat{\xi}^{l}g_{k_{l}}\to\hat{\xi}g$$\hat{\mathbb{P}}$-a.s. By $\hat{\xi}^{l}\stackrel{{\scriptstyle\mathscr{D}}}{{\sim}}\xi$, $l\in\mathbb{N},$ we can conclude $\hat{\xi}\stackrel{{\scriptstyle\mathscr{D}}}{{\sim}}\xi$. Next we show
@@
^$\hat{\mathbb{P}}\Big{(}\|\hat{Y}\|_{\infty}\geq\frac{N_{1}}{2}\Big{)}\leq\frac{\phi}{3},\qquad\hat{\mathbb{P}}\Big{(}\|\hat{Y}^{l}-\hat{Y}\|_{\infty}\geq\max\Big{(}\frac{\eta}{3C_{\mu,\sigma}\bar{K}},\frac{N_{1}}{2}\Big{)}\Big{)}\leq\frac{\phi}{3}$
@@
$\quad\qquad+\int_{0}^{t}|K_{\mu}(s)|\,\mathrm{d}s\Big{(}\sup_{s\in[0,T]}\sup_{x\in[-N_{1},N_{1}]}|\mu_{n_{k_{l}}}(s,x)-\mu(s,x)|+\sup_{s\in[0,T]}|\mu(s,\hat{Y}^{l}_{s})-\mu(s,\hat{Y}_{s}))|\Big{)}$
@@
Note that ^$\hat{\mathbb{P}}(\bar{K}C_{\mu,\sigma}\|\hat{Y}^{l}-\hat{Y}\|_{\infty}\geq\frac{\eta}{3})\leq\frac{\phi}{3}$ and
^$\hat{\mathbb{P}}(\|\hat{Y}\|_{\infty}\vee\|\hat{Y}^{l}\|_{\infty}\geq N_{1})$($\leq\hat{\mathbb{P}}(\{\|\hat{Y}\|_{\infty}\geq N_{1}\}\cup\{\|\hat{Y}^{l}-\hat{Y}\|_{\infty}+\|\hat{Y}\|_{\infty}\geq N_{1}\})$
^($\leq\hat{\mathbb{P}}\Big{(}\Big{\{}\|\hat{Y}\|_{\infty}\geq\frac{N_{1}}{2}\Big{\}}\Big{)}+\hat{\mathbb{P}}\Big{(}\|\hat{Y}^{l}-\hat{Y}\|_{\infty}\geq\frac{N_{1}}{2}\Big{)}$
@@
^$\hat{\mathbb{P}}(\|G^{l}-G\|_{\infty}\geq\eta)$
^($\leq\hat{\mathbb{P}}(\{\|G^{l}-G\|_{\infty}\geq\eta\}\cap\{\|\hat{Y}\|_{\infty}\vee\|\hat{Y}^{l}\|_{\infty}<N_{1}\})+\hat{\mathbb{P}}(\|\hat{Y}\|_{\infty}\vee\|\hat{Y}^{l}\|_{\infty}\geq N_{1})$
@@
Using a uniqueness of limits argument, one obtains $\hat{V}_{t}=\int_{0}^{t}K_{\sigma}(t-s)\,\mathrm{d}\hat{M}_{s}$ for all $t\in[0,T]$ and by the continuity we can conclude that the processes are indistinguishable. Taking a limit in probability in (A.1) or the $\hat{\mathbb{P}}$-a.s. limit of some subsequence, we obtain that $\hat{X}$ is a solution to the SVE (2.1) such that $\mathbb{E}[\sup_{t\in[0,T]}|\hat{X}_{t}-X_{t}|^{2}]\geq\delta$, which contradicts the assumption that pathwise uniqueness holds for SVE (2.1) and, thus, completes the proof. ∎
@@
$\quad+C\int_{0}^{t}\mathbb{E}\big{[}\big{|}X_{s}-\tilde{X}_{s}\big{|}^{p}\big{]}\,\mathrm{d}s\bigg{)}.$

### v3

@@ 1. Introduction
Stochastic Volterra equations (SVEs) are used as mathematical models for the time evolutions of random systems appearing in various areas like biology, finance or physics. SVEs are a natural generalization of ordinary stochastic differential equations (SDEs) and, in contrast to SDEs, they are capable to represent random dynamics with memory effects and very irregular trajectories. For instance, SVEs are used in the modelling of turbulence [BNS08], of volatility on financial markets [EER19] and of deoxyribonucleic acid (DNA) patterns [RBS10].
@@ 2.3. Stability for SVEs
Let $\mu,\mu_{n}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma,\sigma_{n}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$, $n\in\mathbb{N}$, be continuous functions such that:
@@
Suppose that $\|K_{\mu}\|_{q}+\|\tilde{K}_{\mu}\|_{q}<\infty$, $\|K_{\sigma}\|_{2\tilde{q}}+\|\tilde{K}_{\sigma}\|_{2\tilde{q}}$$\infty$, and $\|g\|_{\infty}+\|\tilde{g}\|_{\infty}$$\infty$.
@@ 3.1. Disturbed pendulum equation
Sample paths of the training and the testing sets together with their learned approximations are shown in Figure 1. It is clearly visible that while DeepONet is not able to generalize properly to the testing set, the learned neural SVE paths are very close to the true paths also for the test set.
Neural SVE:
Training set
Neural SVE:
Test set
DeepONet:
Training set
DeepONet:
@@
The results are presented in Table 2. It stands out immediately that for $n=100$ the performance in the test set is far worse than in the training set. Most solutions remain in the range of $\pi,\pi]$. Yet, in some rare cases, solutions may explode outside this range. An example of this can be seen in Figure 2. In the training set this happens too rarely for the neural network to learn the functions outside of $\pi,\pi]$ precisely. Table 2 clearly shows that this discrepancy disappears when the training set is sufficiently large. In the case of $n=2000$ one also sees that the neural network can learn these complex functions as good as in the prior example. Then the plots for the test set look similar to the training set plots of the Neural SVE trained with $n=100$ trajectories.
@@
Neural SVE:
Training set
Neural SVE:
Test set
DeepONet:
Training set
DeepONet:
@@ 3.2. Rough Heston equation
Sample paths of the training and the testing sets together with their learned approximations are shown in Figure 3.
Neural SVE:
Training set
Neural SVE:
Test set
DeepONet:
Training set
DeepONet:
@@ 3.3. Generalized Ornstein–Uhlenbeck process
Sample paths of the training and the testing sets together with their learned approximations are shown in Figure 4.
Neural SVE:
Training set
Neural SVE:
Test set
DeepONet:
Training set
DeepONet:
@@
Sample paths of the training and the testing sets together with their learned approximations are shown in Figure 5.
Neural SVE: Training set Neural SVE: Test set
@@ 3.4. Monetary reserve modelling
As shown in Table 6, the neural SVE learned the dynamics quite well also in this example.
@@
In Figure 6 one can see the log-monetary reserves of the $10$ banks from one data set and its estimation from the test and the training set each. The dips when the first borrowing contracts end after $t=10$ are not as pronounced in the network’s approximation as in the original data set. This likely stems from neural networks struggling to learn indicator functions such as $K_{\mu}$ here. Still, the order of the banks’ reserves as well as their magnitude are captured very well.
@@ 3.5. Comparison to neural SDEs
where $\xi\sim\mathcal{N}(5,0.5)$ and $T=5$. The process $(X_{t})_{t\in[0,5]}$ is expected to decrease in the first quarter of the interval $[0,5]$ where K(ts)=1K(t-s)=1 holds due to the mean-reverting effect of the drift coefficient $\mu(s,x)=2-x$, then something unpredictable will happen and finally in the last part of the interval $t\in[0,5]$ where the kernels attain 1-1 for a large proportion of $s\in[0,t]$, the process might become big due to the turning sign in the drift. Hence, it is expected that the path dependency will have a substantial impact.
@@
Sample paths of the training and the testing sets together with their learned approximations are shown in Figure 7.
Neural SVE:
Training set
Neural SVE:
Test set
Neural SDE:
Training set
Neural SDE:
@@ 3.6. Computational aspects
[…]
@@
[…]
in $C([0,T];\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\times\mathbb{R}^{d})$ as $l\to\infty$, $^\hat{\mathbb{P}}$-a.s., and $\hat{\xi}^{l}\to\hat{\xi}$ as $l\to\infty$, $^\hat{\mathbb{P}}$-a.s. [^1] One may drop the $\xi,\hat{\xi}^{l}$, $l\in\mathbb{N}$, here, as they are uniquely determined by the $X$, $\hat{X}^{l}$, respectively. With $\stackrel{{\scriptstyle\mathscr{D}}}{{\sim}}$ we denote equality in law. From here on we identify any space of continuous functions with the supremum norm.
@@
[…]
@@
We know that $\hat{Y}^{l}\to\hat{Y}$ and that $\hat{\xi}^{l}g_{k_{l}}\to\hat{\xi}g$ $^\hat{\mathbb{P}}$-a.s. By $\hat{\xi}^{l}\stackrel{{\scriptstyle\mathscr{D}}}{{\sim}}\xi$, $l\in\mathbb{N},$ we can conclude $\hat{\xi}\stackrel{{\scriptstyle\mathscr{D}}}{{\sim}}\xi$. Next we show
@@
$\hat{\mathbb{P}}\Big(\|\hat{Y}\|_{\infty}\geq\frac{N_{1}}{2}\Big)\leq\frac{\phi}{3},\qquad\hat{\mathbb{P}}\Big(\|\hat{Y}^{l}-\hat{Y}\|_{\infty}\geq\max\Big(\frac{\eta}{3C_{\mu,\sigma}\bar{K}},\frac{N_{1}}{2}\Big)\Big)\leq\frac{\phi}{3}$
@@
$\quad\qquad+\int_{0}^{t}|K_{\mu}(s)|\,\mathrm{d}s\Big(\sup_{s\in[0,T]}\sup_{x\in[-N_{1},N_{1}]}|\mu_{n_{k_{l}}}(s,x)-\mu(s,x)|+\sup_{s\in[0,T]}|\mu(s,\hat{Y}^{l}_{s})-\mu(s,\hat{Y}_{s})|\Big)$
@@
Note that $\hat{\mathbb{P}}(\bar{K}C_{\mu,\sigma}\|\hat{Y}^{l}-\hat{Y}\|_{\infty}\geq\frac{\eta}{3})\leq\frac{\phi}{3}$ and
$\hat{\mathbb{P}}(\|\hat{Y}\|_{\infty}\vee\|\hat{Y}^{l}\|_{\infty}\geq N_{1})$ $\leq\hat{\mathbb{P}}(\{\|\hat{Y}\|_{\infty}\geq N_{1}\}\cup\{\|\hat{Y}^{l}-\hat{Y}\|_{\infty}+\|\hat{Y}\|_{\infty}\geq N_{1}\})$
$\leq\hat{\mathbb{P}}\Big(\Big\{\|\hat{Y}\|_{\infty}\geq\frac{N_{1}}{2}\Big\}\Big)+\hat{\mathbb{P}}\Big(\|\hat{Y}^{l}-\hat{Y}\|_{\infty}\geq\frac{N_{1}}{2}\Big)$
@@
$\hat{\mathbb{P}}(\|G^{l}-G\|_{\infty}\geq\eta)$
$\leq\hat{\mathbb{P}}(\{\|G^{l}-G\|_{\infty}\geq\eta\}\cap\{\|\hat{Y}\|_{\infty}\vee\|\hat{Y}^{l}\|_{\infty}<N_{1}\})+\hat{\mathbb{P}}(\|\hat{Y}\|_{\infty}\vee\|\hat{Y}^{l}\|_{\infty}\geq N_{1})$
@@
Using a uniqueness of limits argument, one obtains $\hat{V}_{t}=\int_{0}^{t}K_{\sigma}(t-s)\,\mathrm{d}\hat{M}_{s}$ for all $t\in[0,T]$ and by the continuity we can conclude that the processes are indistinguishable. Taking a limit in probability in (A.1) or the $^\hat{\mathbb{P}}$-a.s. limit of some subsequence, we obtain that $\hat{X}$ is a solution to the SVE (2.1) such that $\mathbb{E}[\sup_{t\in[0,T]}|\hat{X}_{t}-X_{t}|^{2}]\geq\delta$, which contradicts the assumption that pathwise uniqueness holds for SVE (2.1) and, thus, completes the proof. ∎
@@
$\quad+C\int_{0}^{t}\mathbb{E}\big[\big|X_{s}-\tilde{X}_{s}\big|^{p}\big]\,\mathrm{d}s.$